Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

a probability problem, who can help? Thanks!

Expert replies
by billzhao » Wed Feb 18, 2009 5:01 pm
A man chooses an outfit from 3 different shirts, 2 different pairs of shoes, and 3 different pants. If he randomly selects 1 shirt, 1 pair of shoes, and 1 pair of pants each morning for 3 days, what is the probability that he wears the same pair of shoes each day, but that no other piece of clothing

Answer is


[/spoiler]
Yiliang
Join the discussion
Source: — Problem Solving |

by earth@work » Wed Feb 18, 2009 6:37 pm
is the OA 1/18; also some part of the question seems to be missing in the end...i guess it says none of the other clothes are repeated

1st day: P1=1/3*1/3 =1/9
2nd day: P2=1/2*1/2 =1/4
3rd day:p3=1*1=1

total P=1/9*1/4*1 = 1/(9*4)
this can be done in 2 ways =2/(9*4)=1/18
Join the discussion

by DarkKnight » Wed Feb 18, 2009 6:58 pm
I am getting 14/81.

(9*8*7*1)*2 / (9*2*9*2*9*2)
Join the discussion

by Bidisha_800 » Thu Feb 19, 2009 12:24 am
total number of combination =3*2*3=18

day 1 = (3 X 1 (shoe #1) X3)/18= 9/18

day2 =(2X 1 (shoe#1) X2)/18 = 4/18

day 3 = 1 X 1 X 1/18 =1/18

same can happen with shoe #2

so probability = 2* (9/18)* (4/18)*(1/18)

=1/81
GMATPowerPrep Test1= 740
GMATPowerPrep Test2= 760
Kaplan Diagnostic Test= 700
Kaplan Test1=600
Kalplan Test2=670
Kalplan Test3=570
Join the discussion

by Mr2Bits » Thu Feb 19, 2009 8:52 am
Join the discussion

by billzhao » Thu Feb 19, 2009 8:43 pm
To solve this, first realize that you don’t need to do calculations for day one; the odds of him wearing any outfit are 1. For day two, he needs to wear one of the two other shirts that he didn’t wear on day one, the same pair of shoes, and one of the two other pants; the probability of this is: 2/3*1/2*2/3=2/9 . For day three, he needs to wear the remaining unworn shirt, the same pair of shoes, and the remaining unworn pants; the probability is:1/3*1/2*1/3=1/18 . So, the probability that he will wear a different shirt and pants on all three days is:2/9*1/18=1/81=(1/3)^4 .
Yiliang
Join the discussion